# MDDM: A Molecular Dynamics Diffusion Model to Predict Particle Self-Assembly

### v2

@@ Abstract
The discovery and study of new material systems relies on molecular simulations that often come with significant computational expense. We propose MDDM, a Molecular Dynamics Diffusion Model, which is capable of predicting a valid output conformation for a given input pair potential function. After training MDDM on a large dataset of molecular dynamics self-assembly results, the proposed model can convert uniform noise into a meaningful output particle structure corresponding to an arbitrary input potential. The model’s architecture has domain-specific properties built-in, such as satisfying periodic boundaries and being invariant to translation. The model significantly outperforms the baseline point-cloud diffusion model for both unconditional and conditional generation tasks.
@@ 1 Introduction
Molecular Dynamics (MD) is a powerful computational tool that lets scientists and engineers study chemical, biological, or material systems at a micro- or nano-scale. In particular, we target a materials science application of molecular self-assembly in which the goal is to model the dynamics and structure of bulk systems containing many particles that interact with one another via a specified potential energy function. By simulating the motion and interaction of particles in a molecular system, material properties can be measured from the resulting equilibrated particle structures.
@@
While MD undoubtedly provides engineers with the capacity to perform high-fidelity material simulations, it is not without its own limitations, namely computational expense. For one, very large systems (i.e. with many particles) are required to emulate the properties of a bulk material as accurately as possible. Crucially, the number of particle-to-particle interactions in a system scales quadratically with the number of particles. While each time step’s computations can be parallelized across multiple processors, the time-evolving nature of an MD simulation is inherently serial, leading to prohibitively lengthy simulation times.
@@
In this work, we propose a denoising diffusion probabilistic model (for short, a diffusion model) that can act as a surrogate model for large-scale MD. A diffusion model starts with noise, and – through an iterative denoising process – generates a sample from the dataset distribution. Most often, diffusion models are image-generators, producing realistic images, often conditioned on a text prompt. In our proposed Molecular Dynamics Diffusion Model (MDDM), rather than creating an image from an input prompt, a set of particle locations is created from an input pair potential function.
@@
We demonstrate both unconditional and conditional conformation generation for particle self-assembly. Our model properly accounts for the periodic boundary conditions present in MD. MDDM generates structures that match the target radial distribution functions qualitatively and quantitatively better than the baseline. By sampling from the MDDM model, the structure of materials systems can be obtained rapidly, with far fewer diffusion model denoising iterations (hundreds) compared to MD time steps (millions), letting an engineer study bulk material systems faster than ever before.
@@
2 Dataset and Task
@@ 2.1 Dataset
We have generated a large dataset of MD results for this problem using LAMMPS [1] because there is not an existing dataset for this task. The simulations consist of 1000 particles each within a $10\times 10\times 10$ box (units are dimensionless ‘LJ’ units). The system is annealed to the target temperature from a temperature 10×10$\times$10 × larger, and the pair potential is an oscillating pair potential given by $U_{\text{OPP}}(r)=r^{-15}+r^{-3}\cos(k(r-1.25)-\phi)$, for parameters $k$ and $\phi$. Therefore, the inputs and outputs of the simulation are summarized as follows:
@@
    Potential frequency $k$, Potential shift $\phi$, Target system temperature $T$
@@
We performed simulations across the parameter ranges: $1.0\leq k\leq 15.0$;  $0.0\leq\phi\leq 6.0$;  $0.01\leq T\leq 0.05$, with 10 values across each input parameter, for a total of 1000 initial simulations. Visualizations of a few of these results are shown in Figure A1 in the Appendix.
2.2 Task and Evaluation Metric
The final goal is to, given the MD inputs from above, generate a set of particle locations that is indistinguishable from an MD trajectory output that had the same inputs. That is, the goal of MDDM will be to sample from the distribution of all possible output MD structures that correspond to a given input potential, temperature, and density. The necessary components are summarized in Fig.1.
@@ 2.2 Task and Evaluation Metric
For an evaluation metric, we wish to compare the simulation structure output with those generated via our model. The radial distribution function (RDF) is a useful function that describes the normalized density of a particle system, with respect to a reference particle. RDFs are often visually compared to ensure no significant mismatch, as in [2]. For a more quantitative result, we use a loss function RDF-MSE, which computes the mean-squared-error between two RDF curves. We will also include RDF visualizations for comparison. This is especially important because visual inspection of particle structures themselves is infeasible, as is clear from Figure A1.
3 Related Work
Predicting the results of MD simulations is a highly sought-after task, often referred to as conformation generation [3]. Spellings et al. [4] characterize self-assembled molecular structures by classifying their phase, leveraging Gaussian Mixture Models and various supervised learning techniques. While effective, they do not address the generative problem associated with creating a valid crystal structure. As for generative models, Arts et al. [5] use a diffusion model approach to predict the free energy coarse-grained MD structures, while Wu et al. [6] generate the steady-state conformation of molecules given their connectivity graphs. These approaches are similar to ours, but fail to solve a large-scale self-assembly generation task, as their simulations contain only a few dozen particles, rather than hundreds/thousands. Some authors attempt the inverse problem, trying to recover a suitable potential for a given output simulation [7, 8, 9]. However, these are iterative methods that require performing simulations in-the-loop. A generative surrogate would significantly accelerate such processes.
While the specific problem we pose has not been tackled using Generative AI, generative denoising diffusion models in the point-cloud domain have indeed been recently explored in a number of areas, such as high-energy physics simulation [10], anatomical reconstruction [11], and protein backbone generation [12]. Point-cloud diffusion models can also generate point clouds of common objects, having been trained on their respective shape datasets [13, 14]. Earlier, GANs were used for the same purpose [15]. These methods make use of point cloud encoders such as PointNet [16]. Other methods such as graph networks [17] or transformer-based approaches [18] can be used for this intermediate task as well.
4 Methods
In this section, we define the proposed MDDM model and outline procedures for training and sampling from the model. We also define a baseline to compare our method against.
Our approach is a denoising diffusion probabilistic model that operates in the 3-D point domain (rather than on 2-D images, as is typical of diffusion models). The input to our denoising process will be a set of particle coordinates sampled at random. These particle locations will be passed into a noise predictor model along with $k,\phi,$ and $T$ (the inputs to the MD system – akin to the text prompt in an image diffusion model). The resulting prediction is a displacement of each particle, which we add to the initial positions to yield the denoised positions. This denoising procedure is repeated $T$ iterations, and the final output should be a particle conformation that corresponds to the inputs.
4.1 Benchmark
Due to of the lack of generative models for self-assembly in the MD domain, we compare against a somewhat disparate baseline: Diffusion with Transolver. That is, we use Transolver [18] as a means of iteratively denoising a point cloud from noise to structure. The node-embedding capabilities of Transolver make it ideal for processing an input shape and predicting a suitable noise vector at each node. We apply the same diffusion strategy for the baseline as we do for our proposed model, which was inspired in part by Diffusion Point Cloud [13].
@@ 4.2 Model Architecture
The diffusion model architecture is defined in analogy with image diffusion methods, as shown in Fig. 2.
Figure 2: Text-conditioned image diffusion models, compared with the proposed potential-conditioned particle diffusion model for MD conformation generation
The denoising step has been designed specifically to operate in the presence of periodic boundary conditions. The full denoising model architecture is shown in Fig. 3 below.
@@
Each PBC-Conv layer involves first creating a k-nearest neighbors graph in the periodic domain. We implement a set of terse, yet still further optimizable, functions in PyTorch [19] to generate this. Once generated for a set of particles, the same graph can be re-used for subsequent PBC-Conv layers, until the denoising step has been applied, modifying particle positions. Then Eqn. 1 is a learnable convolution function defined as follows:
@@
where $\mathbf{f}_{i}^{\prime}$ is the output convolved set of features of particle $i$, $\mathcal{N}(i)$ are the periodic $k$-nearest neighbors of particle $i$, $\left[\mathbf{x}_{j}-\mathbf{x}_{i}\right]_{\circlearrowright}$ refers to the vector from neighbor $\mathbf{x}_{j}$ to source $\mathbf{x}_{i}$ (wrapped across periodic boundaries as needed), $\mathbf{f}_{i}$ is the input pre-convolution set of features at particle $i$, and $\mathbf{g}$ is the global feature vector; $\mathbf{g}$ contains information about the MD system, such as potential parameters and temperature, as well as diffusion process parameters, like diffusion fraction $t/T$. This function is based on the EdgeConv operation from Wang et al. [17], but accounts for periodic boundaries and does not explicitly pass absolute coordinates into the model, making our model translation-invariant (also a property of MD simulations).
Our network consists of 8 PBC-Conv layers with hidden local/global feature counts of 32 at each layer, using a periodic $k$-nearest neighbors graph with $k=32$. Each PBC-Conv layer contains an MLP with 2 hidden layers of 32 neurons each. The output MLP has 2 layers, each with 128 neurons. The sole global input was $t/T$, current diffusion time step divided by the total number of diffusion steps, with $T=500$ for our experiments.
@@ 4.3 Training and Sampling
in which $\mathbf{x}_{0}$ is the data, $\mathbf{x}_{t}$ is the data corrupted by $t$ steps of diffusion, and $\alpha_{t}$ is a constant indicating the magnitude of diffusion from step 00 to step $t$. However, use of this equation is predicated on the notion that at the final diffusion step, pixel values are normally distributed. For our systems, due to the presence of PBCs, particle locations will not obey a normal distribution after diffusion. Rather, they will take on a (very nearly) uniform distribution, since wrapping a normal distribution makes it become approximately uniform. A proof of this statement can be found in the Appendix. Because the final noise is not normally distributed, the standard DDPM variational inference algorithm must be slightly reformulated. To do so, we redefine the forward diffusion process as follows:
$\mathbf{x}_{t}=\mathbf{x}_{0}+\sqrt{\alpha_{t}}\boldsymbol{\epsilon},\quad\text{for}\;\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ (3)
Now, we can continue to use properties of Gaussians, and maximization of the Evidence Lower Bound (ELBO) during training can be re-derived for this revised formulation. Note that the result in Eqn. 3 is also wrapped within the periodic box domain. This gives rise to the training and sampling pseudocode below in Algorithms 1 and 2. A probabilistic derivation is provided in the Appendix.
@@
For our experiments, we use $T=500$ diffusion steps with $\alpha$ on a cosine-schedule $\alpha_{t}=\cos^{2}\left(\frac{\pi}{2}\cdot\frac{x/(T+1)+s}{1+s}\right)$, with $s=0.008$. We train for 800 epochs using an Adam optimizer with learning rate 0.005, which decreases by a factor of 0.95 every 100 epochs.
5 Experiments
We run experiments on two versions of the MD task: unconditional generation and conditional generation. For unconditional generation, we train each model to generate a single randomly-selected MD conformation from our dataset – no condition prompt is fed into the model. The dataset is augmented with rotated/flipped versions of this conformation to avoid memorization of coordinate locations. For conditional generation, we train models to generate valid output conformations given the $k$, $\phi$, and $T$ as additional input conditions, for the range of input parameters described in Sec. 2. We now present results for both models on unconditional and conditional tasks, with RDF-MSE results shown in Tab. 1.
Table 1: Performance, as measured by RDF-MSE, of our method compared to the baseline method.

### v3

@@ Abstract
The discovery and study of new material systems rely on molecular simulations that often come with significant computational expense. We propose MDDM, a Molecular Dynamics Diffusion Model, which is capable of predicting a valid output conformation for a given input pair potential function. After training MDDM on a large dataset of molecular dynamics self-assembly results, the proposed model can convert uniform noise into a meaningful output particle structure corresponding to an arbitrary input potential. The model’s architecture has domain-specific properties built-in, such as satisfying periodic boundaries and being invariant to translation. The model significantly outperforms the baseline point-cloud diffusion model for both unconditional and conditional generation tasks.
@@ 1 Introduction
Molecular Dynamics (MD) is a powerful computational tool that lets scientists and engineers study chemical, biological, or material systems at a micro- or nano-scale. We target a materials science application of molecular self-assembly in which the goal is to model the dynamics and structure of bulk systems containing many particles that interact with one another via a specified potential energy function. By simulating the motion and interaction of particles in a molecular system, material properties can be measured from the resulting equilibrated particle structures.
@@
While MD provides engineers with the capacity to perform high-fidelity material simulations, it is not without its own limitations, namely computational expense. For one, to emulate the properties of a bulk material as accurately as possible, very large systems (i.e. systems with many particles) are required. Crucially, the number of particle-to-particle interactions in a system scales quadratically with the number of particles. While each time step’s computations can be parallelized across multiple processors, the time-evolving nature of an MD simulation is inherently serial, leading to prohibitively lengthy simulation times, especially for many-particle systems.
Predicting the final set of particle locations in an MD simulation, often referred to as conformation generation [1], is a popular area of interest in molecular research, which has been approached from a number of machine learning-based directions. Spellings et al. [2] characterize self-assembled molecular structures by classifying their phase, leveraging Gaussian Mixture Models and various supervised learning techniques. While effective, they do not address the generative problem associated with creating a valid crystal structure. As for generative models, Arts et al. [3] use a diffusion model approach to predict the free energy coarse-grained MD structures, while Wu et al. [4] generate the steady-state conformation of molecules given their connectivity graphs. These approaches are similar to ours, but fail to solve a large-scale self-assembly generation task, as their simulations contain only a few dozen particles, rather than hundreds/thousands. Some authors attempt the inverse problem, trying to recover a suitable potential for a given output simulation [5, 6, 7]. However, these are iterative methods that require performing simulations in-the-loop. A generative surrogate would significantly accelerate such processes. GeoDiff, a diffusion model proposed by Xu et al. [8], features a custom roto-translational invariant graph convolution in a diffusion framework to generate conformations of molecules. Our work is similar, but it differs in that we extend the theory of denoising diffusion such that periodic boundary conditions can be satisfied; we also target large-scale material simulation rather than structure generation for molecules.
While the specific problem we pose has not been tackled using generative AI, generative denoising diffusion models in the point-cloud domain have indeed been recently explored in a number of areas, such as high-energy physics simulation [9], anatomical reconstruction [10], and protein backbone generation [11]. Point-cloud diffusion models can also generate point clouds of common objects, having been trained on their respective shape datasets [12, 13]. Earlier, Generative Adversarial Networks (GANs) were used for the same purpose [14]. These methods make use of point cloud encoders such as PointNet [15]. Other methods such as graph networks [16] or transformer-based approaches [17] can be used for this intermediate task as well.
@@
In this work, we propose a denoising diffusion probabilistic model (i.e., a diffusion model) that can act as a surrogate model for large-scale MD. A diffusion model starts with noise, and – through an iterative denoising process – generates a sample from the dataset distribution. Most often, diffusion models are image-generators, producing realistic images, often conditioned on a text prompt. In our proposed Molecular Dynamics Diffusion Model (MDDM), rather than creating an image from an input prompt, a set of particle locations is created from an input pair potential function. This distinction is illustrated in Fig. 1.
Figure 1: Text-conditioned image diffusion models, compared with the proposed potential-conditioned particle diffusion model for MD self-assembly conformation prediction
@@
We demonstrate both unconditional and conditional conformation generation for particle self-assembly. MDDM properly accounts for the periodic boundary conditions present in MD. Our model generates structures that match the target radial distribution functions qualitatively and quantitatively better than the baseline. By sampling from the MDDM model, the structure of materials systems can be obtained rapidly, with far fewer diffusion model denoising iterations (hundreds) compared to MD time steps (millions), letting an engineer study bulk material systems more efficiently than ever before.
@@
2 Methods
In this section, we formalize the self-assembly conformation generation task we aim to solve using a diffusion-based approach. We also describe the dataset, define the proposed MDDM model, and outline procedures for training and sampling from the model.
@@ 2.1 Dataset
We have generated a large dataset of MD results for this problem using LAMMPS [18] because there is not an existing dataset for this task. The simulations consist of 1000 particles each within a $10\times 10\times 10$ box (units are dimensionless ‘LJ’ units). The system is annealed to the target temperature from a temperature $10\times$ larger, and the pair potential is an oscillating pair potential (OPP) given by $U_{\text{OPP}}(r)=r^{-15}+r^{-3}\cos(k(r-1.25)-\phi)$, for parameters $k$ and $\phi$. This potential was introduced by Mihalkovič and Henley [19] as a concise potential form capable of producing the complex behavior of quasi-crystal structures. The inputs and outputs of the simulation are summarized as follows:
@@
    Potential frequency parameter $k$; Potential shift parameter $\phi$; Target system temperature $T$
@@
We performed simulations across the parameter ranges: $1.0\leq k\leq 15.0$;  $0\leq\phi\leq 6.0;  0.0$$01\leq T\leq 0.05$, with 10 values across each input parameter, for a total of 1000 initial simulations. Visualizations of a few of these results are shown in Figure A1 in the Appendix.
2.2 Task and Benchmark
The final goal is to, given the MD inputs from above, generate a set of particle locations that is thermodynamically indistinguishable from an MD trajectory output that had the same inputs. In probabilistic modeling terms, the goal of MDDM will be to sample from the distribution of all possible output MD structures that correspond to a given input potential and temperature. The necessary components are summarized in Fig. 2, which illustrates the inputs and outputs, as well as the flow of modular scripts we have implemented to perform each step.
@@ 2.2 Task and Benchmark
For an evaluation metric, we wish to compare the simulation structure output with those generated via our model. The radial distribution function (RDF) is a useful function that describes the normalized density of a particle system, with respect to a reference particle. We will also include RDF visualizations for comparison. RDFs are often visually compared to ensure no significant mismatch, as in [20]. This is especially important because visual inspection of particle structures themselves is infeasible, as is clear from Figure A1. For a quantitative result to allow comparison between the results of multiple distinct methods, we use a custom loss function RDF-MSE, which computes the mean-squared-error between two RDF curves. Specifically, the RDF is computed as a normalized histogram of particle-to-particle distances with $n_{b}$ bins spanning from 0 to half of the simulation box side length; the RDF-MSE thus compares two $n_{b}$-element vectors, where $n_{b}=100$ has been selected as a hyperparameter.
Due to of the lack of generative models for material self-assembly in the MD domain, we compare against a somewhat disparate baseline: Diffusion with Transolver. That is, we use Transolver [17] as a means of iteratively denoising a point cloud from noise to structure. The node-embedding capabilities of Transolver allow it to process an input shape and predict a suitable noise vector at each node. We apply the same diffusion strategy for the baseline as we do for our proposed model, which was inspired in part by Diffusion Point Cloud [12].
@@ 2.3 Model Architecture
Our approach is a denoising diffusion probabilistic model that operates in the 3-D point domain (rather than on 2-D images, as is typical of diffusion models). The input to our denoising process will be a set of particle coordinates sampled at random. These particle locations will be passed into a noise predictor model along with $k,\phi,$ and $T$ (the inputs to the MD system – akin to the text prompt in an image diffusion model). The resulting prediction is a displacement of each particle, which we add to the initial positions to yield the denoised positions. This denoising procedure is repeated for $T$ iterations, and the final output is a particle conformation that corresponds to the inputs.
As previously discussed, the diffusion model architecture is defined in analogy with image diffusion methods, as shown in Fig. 1. The denoiser must therefore operate not on image data, but on a point cloud. Furthermore, our denoising step has been designed specifically to operate in the presence of periodic boundary conditions. Figure 3 depicts the full denoising model architecture.
@@
At each PBC-Conv layer, first a k-nearest neighbors graph is constructed, wrapping properly in the periodic domain. We implement a set of simple functions in PyTorch [21] to generate this; code for a simplified implementation is included in Appendix C. Once generated for a set of particles, the same graph can be re-used for subsequent PBC-Conv layers, until the denoising step has been applied, modifying particle positions. PBC-Conv layers are learnable convolution functions defined as follows:
@@
where $\mathbf{f}_{i}^{\prime}$ is the output convolved set of features of particle $i$, $\mathcal{N}(i)$ are the periodic $k$-nearest neighbors of particle $i$, $\left[\mathbf{x}_{j}-\mathbf{x}_{i}\right]_{\circlearrowright}$ refers to the vector from neighbor $\mathbf{x}_{j}$ to source $\mathbf{x}_{i}$ (wrapped across periodic boundaries as needed), $\mathbf{f}_{i}$ is the input pre-convolution set of features at particle $i$, and $\mathbf{g}$ is the global feature vector; $\mathbf{g}$ contains information about the MD system, such as potential parameters and temperature, as well as diffusion process parameters, like diffusion fraction $t/T$. (Note that $||$ denotes concatenation along the feature dimension.) This function is based on the EdgeConv operation from Wang et al. [16], but accounts for periodic boundaries and does not explicitly pass absolute coordinates into the model, making our model translation-invariant (also a property of MD simulations). Rotation-invariant graph convolutions have been proposed for atomic and molecular contexts like this [8, 22], but we do not explicitly provide our network with this property. Rather, we augment our dataset with flipped, rotated, and axis-swapped data conformations, such that the network can implicitly learn the notion of rotation invariance as needed.
Our network consists of 8 PBC-Conv layers with hidden local/global feature counts of 32 at each layer, using a periodic $k$-nearest neighbors graph with $k=32$. Each PBC-Conv layer contains an MLP with 2 hidden layers of 32 neurons each. The output MLP has 2 layers, each with 128 neurons. For unconditional conformation generation, the sole global input was $t/T$, current diffusion time step divided by the total number of diffusion steps, with $T=500$ for our experiments; for conditional generation, temperature and potential parameters (each scaled within the range [1,1][-1,1]) are set as additional global coordinates.
@@ 2.4 Training and Sampling
in which $\mathbf{x}_{0}$ is the data, $\mathbf{x}_{t}$ is the data corrupted by $t$ steps of diffusion, and $\alpha_{t}$ is a constant indicating the magnitude of diffusion from step 0 to step $t$. However, use of this equation is predicated on the notion that at the final diffusion step, pixel values are normally distributed. For our systems, due to the presence of PBCs, particle locations will not obey a normal distribution after diffusion. Rather, they will take on a (very nearly) uniform distribution, since wrapping a normal distribution makes it become approximately uniform. A proof of this statement can be found in Appendix D. Because the final noise is therefore not normally distributed, the standard DDPM variational inference algorithm must be slightly reformulated. To do so, we redefine the forward diffusion process as follows:
$\mathbf{x}_{t}=\text{wrap}\left(\mathbf{x}_{0}+\sqrt{\alpha_{t}}\boldsymbol{\epsilon}\right),\quad\text{for}\;\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}).$ (3)
Now, we can continue to use properties of Gaussian distributions, and maximization of the Evidence Lower Bound (ELBO) during training can be re-derived for this revised formulation. Note that the result in Eqn. 3 is also wrapped within the periodic box domain. This gives rise to the training and sampling pseudocode in Algorithms 1 and 2. A probabilistic derivation of the mean and variance seen in the denoising step of Algorithm 2 is provided in Appendix D. For our experiments, we use $T=500$ diffusion steps with $\alpha$ on a cosine-schedule $\alpha_{t}=\cos^{2}\left(\frac{\pi}{2}\cdot\frac{x/(T+1)+s}{1+s}\right)$, with $s=0.008$. We train for 800 epochs using an Adam optimizer with learning rate 0.005, which decreases by a factor of 0.95 every 100 epochs.
@@
3 Experiments and Results
We run experiments on two versions of the MD task: unconditional generation and conditional generation. For unconditional generation, we train each model to generate a single randomly-selected MD conformation from our dataset – no condition prompt is fed into the model. The conditional generative model, on the other hand, is trained on MD conformations from several potential parameter and temperature combinations, to enable generation of conformations that correspond to particular “prompts” of input temperature and potential parameter values. The datasets are augmented with rotated/flipped versions of this conformation to further impart translation/rotation invariance, and to avoid memorization of coordinate locations.
Training was carried out and performance was evaluated for two models: Diffusion with Transolver and MDDM. The Transolver model, with 128,867 parameters, achieved a training time of 24.03 seconds per 1,000 iterations and a sampling time of 2.53 seconds for 500 steps per data. In comparison, MDDM, with 101,891 PBC-Conv parameters, recorded a slightly faster training time of 23.15 seconds per 1,000 iterations but a longer sampling time of 4.17 seconds for 500 steps per data. These times are for a single NVIDIA RTX A6000 GPU with 128GB DDR4 RAM. The increased sampling time for our method can be partially attributed to the repeated graph generation, which requires calculating each particle’s relative distance to its neighbors once at the start of each denoising step. This added computational complexity likely contributes to the observed difference in sampling efficiency.
For conditional generation, we train models to generate valid output conformations given $k$, $\phi$, and $T$ as additional input conditions, for the range of input parameters described in Sec. 2.1. To evaluate our diffusion model, we pass a set of MD inputs into the model and compare the resulting structures using the mean squared error (MSE) between their radial distribution functions (RDFs). We present results for both models on unconditional and conditional tasks, with average RDF-MSE values shown in Tab. 1.
For the unconditional task, the reverse noising process for sampling is visualized in Fig. 4; by the end of the denoising process, at time step 0, we see a close match between the generated outputs and the ground truth – this, along with the low MSE value in Tab. 1, indicates that the denoising procedure can indeed perform large-scale MD conformation generation. The baseline diffusion with Transolver model, however, struggles to yield performance as strong as our MDDM model, likely due to it prioritizing recreation of global shape information rather than local structure. We further observe that for MDDM, within only the first 10 of 500 steps, the RDF has nearly converged, with only small improvements occurring for subsequent sampling time steps, although the RDF-MSE score continues to decrease, indicating additional improvement. Ideally, there would be a more uniform transformation from the noise distribution to the data distribution. To achieve this in future work, we can explore implicit diffusion modeling [23], alternative noise scheduling approaches [24], or diffusion-adjacent methods like flow-matching [25].
Table 1: RDF-MSE performance of our method compared to the baseline method